If a pine cone falls from the top of a 96-foot pine tree, how long does it take to reach the ground?
Answer by scott8148(6628)

the pinecone starts from rest (no initial velocity), so the distance is equal to the product of
__ one-half the acceleration and the square of the time

d=.5at^2 __ the distance is 96 ft and the acceleration (g) is 32 ft/sec^2

96=.5(32)t^2 __ dividing by 16 __ 6=t^2 __ sqrt(6)=t